### v1

# Vision-Language Meets the Skeleton: Progressively Distillation with Cross-Modal Knowledge for 3D Action Representation Learning

Abstract
Supervised and self-supervised learning are two main training paradigms for skeleton-based human action recognition. However, the former one-hot classification requires labor-intensive predefined action categories annotations, while the latter involves skeleton transformations (e.g., cropping) in the pretext tasks that may impair the skeleton structure. To address these challenges, we introduce a novel skeleton-based training framework (C2VL) based on Cross-modal Contrastive learning that uses the progressive distillation to learn task-agnostic human skeleton action representation from the Vision-Language knowledge prompts. Specifically, we establish the vision-language action concept space through vision-language knowledge prompts generated by pre-trained large multimodal models (LMMs), which enrich the fine-grained details that the skeleton action space lacks. Moreover, we propose the intra-modal self-similarity and inter-modal cross-consistency softened targets in the cross-modal contrastive process to progressively control and guide the degree of pulling vision-language knowledge prompts and corresponding skeletons closer. These soft instance discrimination and self-knowledge distillation strategies contribute to the learning of better skeleton-based action representations from the noisy skeleton-vision-language pairs. During the inference phase, our method requires only the skeleton data as the input for action recognition and no longer for vision-language prompts. Extensive experiments show that our method achieves state-of-the-art results on NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD datasets. The code will be available in the future.
Index Terms:
Action Recognition, Vision-Language, Cross-Modal, Self-Supervised, Contrastive Learning, Large Multimodal Models.
I Introduction
Action recognition has become a hot topic due to its wide range of application fields, such as human-computer interaction, sports analysis, and dangerous behavior warning [1], etc. In recent years, the continuous development of depth sensors like Kinect [2] has facilitated the easy acquisition of the human body’s skeleton (joints coordinate). Compared with RGB-D sequences, the 3D skeleton is naturally robust to light, background noise, and viewpoint changes while also having the advantages of being lightweight and privacy-preserving.
Figure 1: Our proposed approach (C2VL) utilizes the vision-language knowledge prompts as supervision to learn task-agnostic 3D human action representations.
Skeleton-based action recognition methods are primarily categorized into two distinct paradigms: supervised learning and self-supervised learning. In the former, skeleton data serves as the input, with human-designed action category labels represented as one-hot vectors employed for supervision to design complex network architectures [3, 4, 5]. This training process aims to learn task-specific spatial-temporal representations for actions, which has achieved impressive performance. On the other hand, self-supervised learning approaches generate positive-negative pairs by designing pretext tasks of data transformations (rotation, masking, and cropping, etc.) to explore task-agnostic action representation in a label-free manner based on contrastive learning [6, 7, 8]. These methods leverage pretext tasks for unsupervised pre-training and then finetune for downstream tasks. The results suggest that the self-supervised approaches have the potential to surpass supervised learning methods.
However, while the abovementioned methods have proven effective in representing human actions, several challenges remain. (1) Obtaining numerous labeled data is extremely expensive and labor-intensive for supervised learning. Additionally, the reliance on one-hot vector classification restricts the generality of these methods since additional labeled data is needed to address any other tasks. (2) For self-supervised learning, skeleton data transformations (rotation, masking, cropping, etc.) of the pretext task may impair the structural information inherent in skeleton action, leading to information bias. Inspired by the recent success in multi-modal contrastive learning [9, 10, 11], we consider designing a novel skeleton training paradigm guided by vision-language knowledge to obviate numerous label annotations or compromise the integrity of the skeleton structures, as shown in Fig. 1.
Specifically, our approach leverages sample-specific vision-language knowledge prompts as the supervision to guide the skeleton encoder in learning task-agnostic action representation, which is called C2VL. Firstly, we utilize LMMs (Grounding DINO [12] and LLaVA [13]) as the vision-language engines, employing text prompts and visual questions to generate one-to-one vision knowledge prompts related to human action images and one-to-one language knowledge prompts corresponding to action descriptions, respectively. Subsequently, the vision-language encoders extract features from these knowledge prompts to establish the vision-language action concept space, enriching object-related details and fine-grained descriptions absent in the skeleton action space. However, these one-to-one format skeleton-knowledge pairs exhibit many-to-many correspondences among different samples because LMMs generate similar semantic descriptions for different skeleton actions, which causes the noisy skeleton-vision-language joint action space. Therefore, we introduce the intra-modal self-similarity and inter-modal cross-consistency softened targets in the cross-modal self-knowledge distillation learning to progressively control and guide the degree of bringing vision-language knowledge prompts and corresponding skeletons closer in the noisy joint space. This soft contrastive learning process effectively guides the skeleton encoder to learn better task-agnostic 3D human action representation than the hard alignment in CLIP methods. During the inference phase, only skeleton data is needed as the input for action recognition, and vision-language knowledge prompts are omitted without incurring additional computational costs.
Our contributions can be summarized as follows:
  • To the best of our knowledge, this is the first work that utilizes vision-language knowledge prompts as supervision to learn task-agnostic skeleton-based action representations in the training phase. During the inference phase, only the skeleton data is used as the input for skeleton-based action recognition.
  • We employ LMMs as the offline vision-language engines to generate vision-language knowledge prompts concerning action descriptions to establish action concept space, which enriches object-related and fine-grained details that the skeleton action space lacked.
  • We introduce the intra-modal self-similarity and inter-modal cross-consistency softened targets to relax the strict one-to-one constraint, effectively guiding the instance discrimination and knowledge distillation process in the noisy joint action space.
  • Extensive experiments show the superior performance of our training paradigm on the NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD datasets. Notably, our self-supervised method even surpasses some supervised learning methods and has a significant margin improvement under semi-supervised learning scenarios.
II Related Work
In this section, we mainly introduce the related work of skeleton-based action recognition, multi-modal representation learning, and vision-language pretraining with noisy data.
II-A Skeleton-based Action Recognition
Supervised skeleton-based action recognition has been the dominant paradigm. CNN-based approaches [14, 15] transform skeleton data into image-like structures, and RNN-based methods [16, 17] delve into long-term contextual relationships from joint sequences, both of them ignore the skeleton’s inherent graph topology. GCN-based methods [3, 18, 4] regard human joints and bones as graph nodes and edges, respectively, enabling effective exploration of the spatial-temporal representation of actions while keeping structural properties. However, the above-mentioned supervised approaches have the disadvantage of requiring extensive human-designed category annotations. Much research pays attention to self-supervised approaches, which are adept at learning task-agnostic representation in a label-free manner. Reconstruction and contrastive learning are two main problem-solving concepts. They reconstruct skeleton structures by masked modeling [19, 20, 21, 22] or generate positive-negative pairs [8, 7, 6, 23, 24] through skeleton transformation (e.g., rotation, cropping) respectively to explore action representation. Unfortunately, these pretext tasks impair the topological structure of the skeleton, leading to information bias. [Summary]: In this paper, we first explore the possibility of learning action representation for skeleton-based action recognition from other modalities (vision and language) without label annotation or skeleton damage.
II-B Multimodal Representation Learning
Numerous multimodal approaches have demonstrated the capacity to learn powerful representations from different modalities. Inspired by their success, foundational vision-language models such as CLIP [9] and ALIGN [25] have gained widespread adoption in the domain of RGB-based action recognition [26, 1, 27, 10, 28, 29, 30, 31, 11]. Many approaches rely on human-designed coarse-grained category names for video action classification [10], pose estimation [28] and action generation [29], etc. Much recent research gradually employs LLMs (GPT-3, etc.) to generate detailed action descriptions of category names for RGB-based action classification [30] and generation [31]. GAP [11] is the first work to utilize LLMs (GPT-3) for skeleton-based action recognition. Regrettably, although these approaches enrich category names with meaningful semantics, the relationship between each action description and skeleton sample falls under the one-to-many format, which belongs to the supervised learning paradigm. [Summary]: In contrast, our proposed framework differs in two aspects. Firstly, we leverage LMMs (Grounding DINO [12] and LLaVA [13]) as offline engines to generate sample-level vision-language knowledge prompts in a one-to-one format without requiring labor-intensive annotations. Second and more importantly, we utilize these one-to-one vision-language knowledge prompts as the supervisory signal to learn skeleton representation in a self-supervised learning paradigm.
II-C Contrastive Learning with Noisy Pairs
Contrastive learning is the core of the recent multimodal self-supervised learning methods, which pulls the positive pairs closer and pushes the negative pairs away [32, 33]. However, the noisy many-to-many correspondences existed in the label-intensive annotated and web-collected pairs, showing the potential semantic similarity between negative samples. This phenomenon leads to the computational inefficiency of the conventional contrast process that pursues the hard alignment [34]. Therefore, several techniques have been developed from the soft alignment perspective to learn robust representations from noisy pairs [35, 36, 34, 37, 38, 39], including the weighting mechanism [35, 36], k-nearest neighbors scheme [39], and knowledge distillation as softening targets [34, 37, 38]. [Summary]: In this paper, we design two softened targets with the self-distillation mechanism to progressively mitigate the adverse effects of noisy pairs and refine inconsistent alignment for more coherent representations.
III Method
Figure 2: The pipeline of our proposed approach. Before the pre-training phase, the vision and language knowledge prompts are generated regarding skeleton sequences by offline LMMs (Grounding DINO [12] and LLaVA [13]) with text prompts and visual questions. In the pre-training phase, the skeleton data is utilized as the input for the skeleton encoder to learn action representation in the skeleton action space. The vision encoder and language encoder are employed to extract features from vision and language knowledge prompts, contributing to the creation of the vision-language action concept space that enhances fine-grained details not captured in the skeleton space. Subsequently, the degree to which pairs consisting of vision-language knowledge prompts and their corresponding skeleton are brought closer should be progressively guided by the intra-modal self-similarity and inter-modal cross-consistency softened targets. During the inference phase, we only utilize the former pre-trained skeleton encoder with a fully connected layer and skeleton data for skeleton-based action recognition without vision-language knowledge prompts.
In this section, we begin with an introduction to generating vision-language knowledge prompts (Section III-A). Following that, we describe the contrastive learning process with InfoNCE loss (Section III-B). Then, we introduce two soft targets (Section III-C). Finally, the progressive training objectives with dynamic mechanisms are described (Section III-D). The pipeline of the method is shown in Fig. 2.
III-A Vision-Language Action Concept Space
For a given skeleton dataset $D\in\left\{S_{i}\right\}_{i=1}^{\left|\mathrm{D}\right|}$, we first generate the sample-level vision-language knowledge prompts $\left\{(V_{i},L_{i})\right\}_{i=1}^{\left|\mathrm{D}\right|}$ by offline LMMs (Grounding DINO [12] and LLaVA [13]). These sample-level vision-language knowledge prompts have a one-to-one corresponding with skeleton sequence, which establishes the vision-language action concept space as shown in Fig. 3.
III-A1 Vision Knowledge Prompts
To the best of our knowledge, prior skeleton-based action recognition methods have yet to explore the possibility of utilizing videos to guide the learning of skeleton representation. Here, we introduce the video modality to enrich the motion patterns and object-related details of skeleton representation. These raw videos and skeleton sequences have one-to-one correspondence as they are captured simultaneously by Kinect devices. However, raw videos are challenged to capture fine-grained details and subtle movement changes because of interference from irrelevant background environments. To mitigate this issue, we leverage a pre-trained open-set object detector called Grounding DINO [12], with text prompts “person” to crop the human action images as vision knowledge prompts precisely. These vision knowledge prompts only consist of human action regions enhancing the motion patterns in the vision action concept space.
III-A2 Language Knowledge Prompts
Inspired by the success of CLIP in handling image-text pairs, numerous methods [10, 28, 29, 11] employ human-designed action names and LLMs-generated movement descriptions for action recognition. However, this language information is at the category level and has a one-to-many relationship with the skeleton sequences, so more sample-specific detail descriptions are needed. Therefore, we aim to establish sample-level language action descriptions to enrich the action concept space without requiring expensive annotation and labor-intensive effort. Fortunately, LMMs have an advantage in visual question answering, presenting a new opportunity. Here, we utilize LLaVA [13] with suitable question prompts “Is he/she or are they holding anything in the hand? Is he/she or are they standing or sitting? What is he/she or are they trying to do? Answer the questions concisely” to comprehend and generate action descriptions to the vision action images autonomously. These language knowledge prompts provide sample-specific fine-grained detail descriptions in the language action concept space.
Figure 3: The establishment of vision-language action concept space.
III-B Contrastive Learning with InfoNCE Loss
Consider a batch of $B$ paired skeleton-vision-language tuples $\left\{(S_{i},V_{i},L_{i})\right\}_{i=1}^{B}$ drawn from the above-mentioned multi-modal space, we aim to learn skeleton encoders $E_{s}$ based on cross-modal contrastive learning. Specifically, for the $i$-th pair, the skeleton data $S_{i}$, vision data $V_{i}$ and language data $L_{i}$ are input into the skeleton encoder $E_{s}$, vision encoder $E_{v}$ and language encoder $E_{t}$ respectively to get the corresponding modality features. Then, various projectors map these features to a shared latent space for generating L2-normalized embedding pairs $\left\{(\mathrm{\mathbf{s}}_{i},\mathrm{\mathbf{v}}_{i},\mathrm{\mathbf{l}}_{i})\right\}_{i=1}^{\mathrm{B}}$, where $\mathrm{\mathbf{s}}_{i},\mathrm{\mathbf{v}}_{i},\mathrm{\mathbf{l}}_{i}\in\mathbb{R}^{d}$. We use the InfoNCE [40] loss to align the skeleton and vision spaces, which focuses on maximizing the similarity between $\mathrm{\mathbf{s}}_{i}$ and $\mathrm{\mathbf{v}}_{i}$ to pulls the paired skeleton-vision embeddings together in two directions. The loss $\mathcal{L}_{s2v}$ for aligning skeleton to vision is defined as follows:
$\mathcal{L}_{s2v}=-\frac{1}{\mathrm{B}}\sum\limits_{i=1}^{\mathrm{B}}\mathrm{log}\frac{\mathrm{exp}(\mathrm{sim}(\mathrm{\mathbf{s}}_{i},\mathrm{\mathbf{v}}_{i})/\tau)}{\sum\nolimits_{j=1}^{\mathrm{B}}\mathrm{exp}(\mathrm{sim}(\mathrm{\mathbf{s}}_{i},\mathrm{\mathbf{v}}_{j})/\tau)},$ (1)
where $\mathrm{sim}(\cdot)$ denotes the cosine similarity, and $\tau$ is the learnable temperature parameter. The alignment process between these two spaces can be denoted as $\mathcal{L}_{\mathrm{infoNCE}}^{sv}=\mathcal{L}_{s2v}+\mathcal{L}_{v2s}$, and the definition of $\mathcal{L}_{v2s}$ is similar to the $\mathcal{L}_{s2v}$. For convenience, we simplify and re-write the loss in the matrix as:
$\mathcal{L}_{\mathrm{infoNCE}}^{sv}=\mathcal{H}(\mathrm{\mathbf{I_{B}}},\rho(\mathrm{\mathbf{SV}^{T}}))+\mathcal{H}(\mathrm{\mathbf{I_{B}}},\rho(\mathrm{\mathbf{VS}^{T}})),$ (2)
$\mathcal{L}_{\mathrm{infoNCE}}^{sl}=\mathcal{H}(\mathrm{\mathbf{I_{B}}},\rho(\mathrm{\mathbf{SL}^{T}}))+\mathcal{H}(\mathrm{\mathbf{I_{B}}},\rho(\mathrm{\mathbf{LS}^{T}})),$ (3)
where $\mathcal{H}$ is the cross-entropy function, and $\rho$ is the standard softmax function. The $\mathrm{\mathbf{I_{B}}}$ is the identity matrix, and $\mathrm{\mathbf{S}},\mathrm{\mathbf{V}},\mathrm{\mathbf{L}}\in\mathbb{R}^{\mathrm{B}\times d}$ are the matrices that contain a batch of skeleton, vision and language embeddings. Besides, we align the skeleton and language space similarly by minimizing the $\mathcal{L}_{\mathrm{infoNCE}}^{sl}$.
In practice, the strict assumption in InfoNCE loss to bring $(s_{i},v_{i},l_{i})$ closer and push $(s_{i},v_{j},l_{k})$ away is unreasonable for two reasons. First, as shown in Fig. 4, the ground truth vision-language knowledge prompts $(v_{i},l_{i})$ of the given skeleton $s_{i}$ may be incorrect because of hallucination in LMM and poor pose estimation, showing low semantic similarity scores among positive pairs. Second, as shown in Fig. 5, the given skeleton $s_{i}$ would be aligned to unpaired vision-language knowledge prompts $(v_{j},l_{k})$ with shared semantics to different degrees within a larger batch size, resulting in high similarity semantic scores among negative pairs. For these, the three modality pairs exhibit many-to-many correspondences among different samples in a batch. The hard alignment between these spaces based on InfoNCE loss will result in poor skeleton representations, as shown in Fig. 7a and Fig. 7c.
III-C Cross-modal Alignment with Soft Targets
To address the limitations of using standard InfoNCE loss to learn skeleton representation on noisy joint action concept space, we attempt to design two soft targets to guide the learning of cross-modal alignment. These soft targets play crucial roles in repairing a sample with more semantically similar correspondences in other modalities, which relaxes the explicit one-to-one correspondences and improves the implicit many-to-many relationships.
III-C1 Intra-modal Self-similarity Guidance
To estimate the many-to-many relationships within $(s_{i},v_{j},l_{k})$, we propose to utilize the intra-modal self-similarity to guide the learning of skeleton representation from the vision-language knowledge prompts. Specifically, we calculate the skeleton-to-skeleton similarity $\mathrm{\mathbf{SS}^{T}}$, vision-to-vision similarity $\mathrm{\mathbf{VV}^{T}}$, and language-to-language similarity $\mathrm{\mathbf{LL}^{T}}$ in their respective modality space as soft labels. For the training stability, we use the weighted average strategy [37] to fuse the hard and soft labels as the softened targets to supervise the skeleton-to-vision, vision-to-skeleton, skeleton-to-language, and language-to-skeleton correspondences, respectively. These intra-modal self-similarity softened targets within a batch are defined as follows:
$\mathbf{P}^{v2s}=\beta\rho(\mathrm{\mathbf{VV}^{T}})+(1-\beta)\mathrm{\mathbf{I_{B}}},$ (4)
$\mathbf{P}^{l2s}=\beta\rho(\mathrm{\mathbf{LL}^{T}})+(1-\beta)\mathrm{\mathbf{I_{B}}},$ (5)
$\mathbf{P}^{s2v},\mathbf{P}^{s2l}=\beta\rho(\mathrm{\mathbf{SS}^{T}})+(1-\beta)\mathrm{\mathbf{I_{B}}},$ (6)
where $\beta$ is a trade-off hyper-parameter set to 0.2. After replacing the hard targets $\mathrm{\mathbf{I_{B}}}$ with softened targets $\mathbf{P}_{\mathrm{intra}}^{sv}$, $\mathbf{P}_{\mathrm{intra}}^{vs}$, $\mathbf{P}_{\mathrm{intra}}^{sl}$ and $\mathbf{P}_{\mathrm{intra}}^{ls}$, the model can control the degree of attractive and repulsive forces between the skeleton and vision-language embeddings based on their intra-modal self-similarity for more generalized skeleton representation.
III-C2 Inter-modal Cross-consistency Boosting
Although the intra-modal self-similarity relaxes the one-to-one constraint and pursues the many-to-many correspondences among the three modalities, the confidence value of the original positive pairs is still at the domination position. Especially for those faulty positive pairs, the higher confidence may restrain and submerge the process of repairing samples with more semantically similar correspondences. To mitigate this issue, we adopt the cycle consistent prediction strategy [36] to generate the inter-nodal cross-consistency soft targets as defined as follows:
$\mathbf{Q}^{v2s}=\rho(\mathrm{\mathbf{SV}^{T}}+\mathrm{\mathbf{I_{B}}}\mathrm{\mathbf{VS}^{T}}\mathrm{\mathbf{J}}+\mathrm{\mathbf{J}}\mathrm{\mathbf{S}^{T}}\mathrm{\mathbf{V}}\mathrm{\mathbf{I_{B}}}),$ (7)
$\mathbf{Q}^{s2v}=\rho(\mathrm{\mathbf{VS}^{T}}+\mathrm{\mathbf{I_{B}}}\mathrm{\mathbf{SV}^{T}}\mathrm{\mathbf{J}}+\mathrm{\mathbf{J}}\mathrm{\mathbf{V}^{T}}\mathrm{\mathbf{S}}\mathrm{\mathbf{I_{B}}}),$ (8)
$\mathbf{Q}^{l2s}=\rho(\mathrm{\mathbf{SL}^{T}}+\mathrm{\mathbf{I_{B}}}\mathrm{\mathbf{LS}^{T}}\mathrm{\mathbf{J}}+\mathrm{\mathbf{J}}\mathrm{\mathbf{S}^{T}}\mathrm{\mathbf{L}}\mathrm{\mathbf{I_{B}}}),$ (9)
$\mathbf{Q}^{s2l}=\rho(\mathrm{\mathbf{LS}^{T}}+\mathrm{\mathbf{I_{B}}}\mathrm{\mathbf{SL}^{T}}\mathrm{\mathbf{J}}+\mathrm{\mathbf{J}}\mathrm{\mathbf{L}^{T}}\mathrm{\mathbf{S}}\mathrm{\mathbf{I_{B}}}),$ (10)
where $\mathrm{\mathbf{J}}$ is an all-one matrix. The first term of Equation 7-10 aggregates the information from its cross opposite modality for dual-direction alignment, i.e., alignment from vision $\mathrm{\mathbf{v}}_{i}$ to skeleton $\mathrm{\mathbf{s}}_{j}$ is based on the probability that skeleton $\mathrm{\mathbf{s}}_{i}$ matched with vision $\mathrm{\mathbf{v}}_{j}$. Moreover, the second and third terms are designed to avoid positive pairs that are not sound correspondences for boosting negative pairs with high semantic similarity in cross-modalities.
Figure 4: Histogram of similarity scores for positive pairs between skeleton and vision-language knowledge representations in cross-modal space after training with original infoNCE loss.
Figure 5: Histogram of similarity scores for negative pairs between skeleton and vision-language knowledge representations in cross-modal space after training with original infoNCE loss. The red ellipse highlights the movement of arms.
III-D Progressive Self-distillation Training Objectives
Here, we introduce a progressive self-knowledge distillation framework[41] where the two soft targets are dynamically updated at each epoch by the teacher network to guide the learning of the student network. Compared to the conventional knowledge distillation methods utilized the static and separate pre-trained teacher network, the main difference of this self-distillation strategy is that it regards the same model architecture as the student and teacher, which has the advantage of reducing computation and memory costs. Meanwhile, as the training goes on, the well-trained teacher can provide more confidence in learning soft labels for students, which mitigates the adverse effects of noisy pairs and enhances the generalization capability.
To align skeleton space with vision-knowledge spaces, we adopt a progressive scheme [34] to gradually control and balance the roles of two soft targets throughout training. Specifically, we partition a batch of pairs into two subsets with the refreshed hyper-parameter $\alpha\in[0,1]$ at each epoch. The intra-modal self-similarity soft target is utilized to repair the first subset data $\mathrm{B_{intra}}=\left\lfloor\alpha\mathrm{B}\right\rfloor$ for estimation of the many-to-many relationships, while the inter-class modal cross-consistency is employed to guide the second subset data $\mathrm{B_{inter}}=\mathrm{B}-\left\lfloor\alpha\mathrm{B}\right\rfloor$ for boosting negative pairs with high semantic similarity. At the initial training phase, the value of $\alpha$ should be large for training stability. With the increased training iterations, we gradually decrease the value of $\alpha$ by cosine annealing schedule [42] to increase the contribution of alignment extent between skeleton and high semantic similarity vision-language knowledge prompts. The start value is 0.9, and the end value is 0.1. In this way, the skeleton representation can be better learned in the noisy cross-modal pairs.
Altogether, the final training objective can be formulated as follows:
$\mathcal{L}_{\mathrm{Soft}}=\mathcal{L}_{\mathrm{Soft}}^{sv}+\mathcal{L}_{\mathrm{Soft}}^{sl},$ (11)
$\begin{split}\mathcal{L}_{\mathrm{Soft}}^{sv}=\alpha&[\mathcal{H}(\mathbf{P}_{\mathrm{intra}}^{s2v},\rho(\mathrm{\mathbf{\overline{S}}\mathbf{V}^{T}}))+\mathcal{H}(\mathbf{P}_{\mathrm{intra}}^{v2s},\rho(\mathrm{\mathbf{\overline{V}}\mathbf{S}^{T}})]+\\ (1-\alpha)&[\mathcal{H}(\mathbf{Q}_{\mathrm{inter}}^{s2v},\rho(\mathrm{\mathbf{\underline{S}}\mathbf{V}^{T}}))+\mathcal{H}(\mathbf{Q}_{\mathrm{inter}}^{v2s},\rho(\mathrm{\mathbf{\underline{V}}\mathbf{S}^{T}})],\end{split}$ (12)
$\begin{split}\mathcal{L}_{\mathrm{Soft}}^{sl}=\alpha&[\mathcal{H}(\mathbf{P}_{\mathrm{intra}}^{s2l},\rho(\mathrm{\mathbf{\overline{S}}\mathbf{L}^{T}}))+\mathcal{H}(\mathbf{P}_{\mathrm{intra}}^{l2s},\rho(\mathrm{\mathbf{\overline{L}}\mathbf{S}^{T}})]+\\ (1-\alpha)&[\mathcal{H}(\mathbf{Q}_{\mathrm{inter}}^{s2l},\rho(\mathrm{\mathbf{\underline{S}}\mathbf{L}^{T}}))+\mathcal{H}(\mathbf{Q}_{\mathrm{inter}}^{l2s},\rho(\mathrm{\mathbf{\underline{L}}\mathbf{S}^{T}})],\end{split}$ (13)
where $\mathbf{P}_{\mathrm{intra}}^{*}$ is the first $\mathrm{B_{intra}}$ rows of intra-modal self-similarity soft targets, and the $\mathbf{Q}_{\mathrm{inter}}^{*}$ is the last $\mathrm{B_{inter}}$ rows of inter-modal cross-consistency soft targets. $\overline{S},\overline{V},\overline{L}\in\mathbb{R}^{\mathrm{B_{intra}}\times d}$ denotes the first $\mathrm{B_{intra}}$ rows of skeleton, vision, and language embeddings, respectively. Meanwhile, $\underline{S},\underline{V},\underline{L}\in\mathbb{R}^{\mathrm{B_{inter}}\times d}$ denotes the last $\mathrm{B_{inter}}$ rows of respective embeddings.
In the inference phase, only the skeleton sequences and encoder $E_{s}$ are used for skeleton-based action recognition. The vision-language prompts and respective encoders will be dropped.
IV Experiments
IV-A Datasets
IV-A1 NTU RGB+D 60 Dataset [43]
It comprises 60 predefined action categories and 56,880 action samples, containing multiple modalities, including RGB, depth map, skeleton, and infrared (IR) video. The skeleton sequences contain a maximum of two human skeletons, represented by the 3D coordinates of 25 joints. The RGB video has a resolution of $1920\times 1080$. Two benchmarks are provided in this dataset: cross-subject (Xsub) and cross-view (Xview). In Xsub, 20 subjects are used for training, while the remaining subjects are used for testing. In Xview, sequences from views 2 and 3 are designated for training, and the sequences from view 1 are used for testing.
IV-A2 NTU RGB+D 120 Dataset [44]
It is an extension of NTU RGB+D 60 with 120 predefined action categories and 114,480 action samples, which have the same modality as NTU RGB+D 60. Similarly, this dataset provides two benchmarks: cross-subject (Xsub) and cross-setup (Xset). In Xsub, sequences from 53 subjects are utilized for training, while the remaining subjects are reserved for testing. In Xset, sequences from even camera IDs are designated for training, and the rest are employed for testing.
IV-A3 PKU-MMD Dataset [45]
It comprises 51 action categories and almost 20000 action samples with the same multi-modalities as the NTU series datasets. The dataset also provides two benchmarks: cross-subject (Xsub) and cross-view (Xview). In Xsub, sequences from 57 subjects belong to the training set and 9 subjects for the testing. In Xview, sequences from the middle and right views are utilized for training, and the left is used for testing. In this experiment, we adopt the cross-subject benchmark following the previous studies.
TABLE I: Comparison to the state-of-the-art methods for action recognition on the NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD II datasets under the linear evaluation protocol. The best and second-best results are highlighted in bold and underlined, respectively. The J refers to the joint modality, the M denotes the motion modality, and the B is the bone modality.
Method Publication Architecture Modality NTU60 NTU120 PKU-MMD II
xsub xview xsub xset xsub
Single-stream:
AimCLR [39] AAAI’22 GCN J 74.3 79.7 63.4 63.4 -
CPM [46] ECCV’22 GCN J 78.7 84.9 68.7 69.6 48.3
CMD [47] ECCV’22 GRU J 79.8 86.9 70.3 71.5 43.0
HaLP [48] CVPR’23 GRU J 79.7 86.8 71.1 72.2 43.5
ActCLR [8] CVPR’23 GCN J 80.9 86.7 69.0 70.5 -
DMMG [49] TIP’23 GCN J 82.1 87.1 69.6 70.1 -
CSTCN [23] TMM’23 GRU J 83.1 88.7 72.5 77.4 48.0
UmURL [50] ACM MM’23 Transformer J 82.3 89.8 73.5 74.3 52.1
PCM3 [51] ACM MM’23 GRU J 83.9 90.4 76.5 77.5 51.5
KTCL [24] TMM’24 Transformer J 82.4 89.4 74.4 74.5 -
C2VL (Ours) This work GCN J 84.4 89.8 76.0 78.7 52.6
Multi-stream:
3s-CrosSCLR [6] CVPR’21 GCN J+M+B 77.8 83.4 67.9 66.7 21.2
3s-AimCLR [39] AAAI’22 GCN J+M+B 78.9 83.8 68.2 68.8 39.5
3s-CMD [47] ECCV’22 GRU J+M+B 84.1 90.9 74.7 76.1 52.6
3s-CPM [46] ECCV’22 GCN J+M+B 83.2 87.0 73.0 74.0 51.5
3s-ActCLR [8] CVPR’23 GCN J+M+B 84.2 88.8 74.3 75.7 -
2s-DMMG [49] TIP’23 GCN J+M 84.2 89.3 72.7 72.4 -
3s-CSTCN [23] TMM’23 GRU J+M+B 85.8 92.0 77.5 78.5 53.9
3s-UmURL [50] ACM MM’23 Transformer J+M+B 84.4 91.4 75.9 77.2 54.3
3s-PCM3 [51] ACM MM’23 GRU J+M+B 87.4 93.1 80.0 81.2 58.2
3s-C2VL (Ours) This work GCN J+M+B 88.3 92.8 82.5 84.3 60.0
IV-B Implementation Details
All experiments are conducted using the PyTorch framework on an NVIDIA A100 GPU. Each sample is down-sampled to 64 frames using the data pre-processing code from [11]. We use ST-GCN [52] as our skeleton GCN encoder. For the image-text encoder, we utilize the pre-trained ViT-L/14@336px model from CLIP [9] and freeze its parameters during training. The temperature parameter $\tau$ is 0.07. In the pre-training phase, we train the model with the SGD optimizer for 150 epochs with a batch size 400. The initial learning rate is 0.1 and reduced by a factor of 0.1 at epochs 130 and 140. Weight decay is set to 5e-4 following the strategy in [11]. In the inference stage, we utilize multiple evaluation protocols, including the linear evaluation protocol, finetune protocol, semi-supervised evaluation protocol(1%, 5% or 10% labeled data), KNN evaluation protocol, and transfer learning evaluation protocol for action recognition.
IV-C Comparison to the State-of-the-art
To comprehensively evaluate the performance of our method, we conduct comparisons with other state-of-the-art methods across various settings.
IV-C1 Linear Evaluation
The linear evaluation protocol applies a fully connected layer after the parameter-fixed ST-GCN to classify. As shown in Table I, we found that both single-stream and three-stream performance can achieve the best or the second-best results on NTU RGB+D 60 & 120 and PKU-MMD II datasets, indicating that vision-language knowledge prompts have a solid capacity to guide GCN for generalization representation.
IV-C2 Finetune Evaluation
We pre-train ST-GCN and finetune the entire network for action recognition, applying a fully connected layer. Table II compares the results obtained through our pre-training and other training paradigms. Notably, our training paradigm performs better in capturing action patterns and details from vision-language knowledge prompts, outperforming some supervised learning methods (Shift-GCN) and even surpassing state-of-the-art self-supervised methods.
TABLE II: Comparison of action recognition results employing different training paradigms (supervised and self-supervised learning) on the NTU dataset. Supervised learning utilizes one-hot labels for training models, while self-supervised learning utilizes the pretext task to pre-train models and fine-tune them with labels. The best and second-best results are highlighted in bold and underlined, respectively. SL: Supervised Learning, SSL: Self-supervised Learning.
Method Paradigm NTU60 NTU120
xsub xview xsub xset
Single-stream:
ST-GCN [52] SL 81.5 88.3 70.7 73.2
Shift-GCN [53] SL 87.8 95.1 80.9 83.2
[…]
AimCLR [39] SSL 83.0 89.2 77.2 76.1
CPM [46] SSL 84.8 91.1 78.4 78.9
ActCLR [8] SSL 85.8 91.2 79.4 80.9
C2VL (Ours) SSL 88.9 93.2 83.4 84.9
Multi-stream:
3s-ST-GCN [52] SL 85.2 91.4 77.2 77.1
4s-Shift-GCN [53] SL 90.7 96.5 85.9 87.6
[…]
3s-AimCLR [39] SSL 86.9 92.8 80.1 80.9
3s-ActCLR [8] SSL 88.2 93.9 82.1 84.6
3s-C2VL (Ours) SSL 91.8 95.7 87.8 89.2
IV-C3 Semi-supervised Evaluation
Our approach involves pre-training with the entire training data and finetuning the classifier with only 1%, 5%, and 10% of the labeled data, respectively. Table III demonstrates that our method significantly renews the accuracy of the state-of-the-art methods, indicating its effectiveness and strong generalization capabilities in learning high-quality task-agnostic representations without relying on extensive labels.
TABLE III: Comparison of action recognition results under semi-supervised evaluation protocol on NTU RGB+D 60 dataset. The best and second-best results are highlighted in bold and underlined, respectively.
Method xsub xview
1% 5% 10% 1% 5% 10%
ASSL [54] - 57.3 64.3 - 63.6 69.8
MS2L [55] 33.1 - 65.1 - - -
MCC [56] - 47.4 60.8 - 53.3 65.8
ISC [57] 35.7 59.6 65.9 38.1 65.7 72.5
Colorization [19] 48.3 65.7 71.7 52.5 70.3 78.9
Hi-TRS [58] 39.1 63.3 70.7 42.9 68.3 74.8
CPM [46] 56.7 - 73.0 57.5 - 77.1
CMD [47] 50.6 71.0 75.4 53.0 75.3 80.2
HiCo [59] 54.4 - 73.0 54.8 - 78.3
PCM3 [51] 53.8 - 77.1 53.1 - 82.8
UmURL [50] 58.1 72.5 - 58.3 76.8 -
C2VL (Ours) 69.3 79.4 81.8 69.1 82.1 85.5
IV-C4 KNN Evaluation
In this experiment, we utilize the K-nearest neighbors (KNN) as a classifier for action retrieval based on the obtained skeleton representation from the pre-training model. The results on NTU60 and NTU120 datasets are shown in Table IV. Our method achieves the best results among all tasks, surpassing the previous methods by a large margin. Moreover, it demonstrates that the obtained skeleton representation is more discriminative than other methods.
TABLE IV: Comparison to the state-of-the-art methods for action retrieval with the joint stream on NTU RGB+D 60 and NTU RGB+D 120 datasets. The best and second-best results are highlighted in bold and underlined, respectively.
Method NTU60 NTU120
xsub xview xsub xset
LongT GAN [60] 39.1 48.1 31.5 35.5
P&C [61] 50.7 76.3 39.5 41.8
ISC [57] 62.5 82.6 50.6 52.3
AimCLR [39] 62.0 71.5 - -
CMD [47] 70.6 85.4 58.3 60.9
HiCLR [62] 67.3 75.3 - -
HiCo [59] 68.3 84.8 56.6 59.1
HaLP [48] 65.8 83.6 55.8 59.0
UmURL [50] 71.3 88.3 58.5 60.9
PCM3 [51] 73.7 88.8 63.1 66.8
C2VL (Ours) 78.0 88.8 64.0 68.8
IV-C5 Transfer Learning Evaluation
We evaluate the transfer generalizability of the obtained skeleton representation in this part. Specifically, we use the model pre-trained on the NTU60 and NTU120 datasets to finetune the PKU-MMD II dataset. As shown in Table V, our method outperforms the competitors by a large margin, showing the excellent transferability of our learned skeleton representation, which has enormous potential in data-scarcity scenarios.
TABLE V: Comparison to the state-of-the-art methods with transfer learning on xsub task. The source dataset is NTU RGB+D 60 or NTU RGB+D 120, and the target dataset is PKU-MMD II. The best and second-best results are highlighted in bold and underlined, respectively.
Method Transfer to PKU-MMD II
NTU60 NTU120
LongT GAN [60] 44.8 -
M2L [55] 45.8 -
ISC [57] 45.9 -
CrosSCLR [6] 54.0 52.8
CMD [47] 56.0 57.0
HiCo [59] 56.3 55.4
UmURL [50] 58.2 57.6
C2VL (Ours) 67.5 69.6
IV-D Evaluation of Knowledge Necessity
To further explore the significance of distinct modal information during the pre-training phase, we employed single-modal knowledge prompts as the supervision for pre-training. Table VI presents the action recognition accuracy of linear evaluation protocol across different modalities. Notably, vision-language knowledge prompts exhibit the best performance and offer valuable guidance and assistance for skeleton representation learning. In addition, language knowledge prompts tend to offer more fine-grained and local semantic details than vision knowledge prompts cannot provide.
TABLE VI: Analysis of different knowledge prompts on NTU RGB+D 60 xsub dataset with the joint stream.
Knowledge xsub xview
Vision Language
78.6 81.9
83.8 88.6
84.4 89.8
IV-E Ablation Study
IV-E1 Influences of Components
We evaluate the efficiency of different components during the pre-training phase in Table VII. It is noted that the proposed soft targets are necessary as they can draw the unpaired data with high similarity closer, fostering more robust task-agnostic skeleton representations. Additionally, dynamic partitioning and progressive mechanisms have effectively controlled the degree of pair-bringing.
TABLE VII: Analysis of different components on NTU RGB+D 60 xsub dataset with the joint stream.
Soft Targets Contrastive Strategies Acc (%)
Intra- Similarity Inter- Consistency Dynamic Partitioning Progressive Control
82.0
82.8
83.0
83.4
[…]
84.0
84.4
IV-E2 Influences of Skeleton Encoders
Furthermore, we analyze the performance of our method across different skeleton encoders, including GRU, GCN, and Transformer architectures. As shown in Table VIII, the GCN series encoders achieve the promised performance, which retains the topology structure of skeleton data and explores the spatial-temporal characteristic of actions. It is noted that the performance positively correlates with the ability of encoders. However, the results of UmURL drop a lot because its Transformer architecture decouples and impairs the spatial-temporal topology property of the skeleton. The PCM3 with GRU architecture neglects the spatial structure of the skeleton, resulting in inferior modeling of spatial movements. Therefore, our method can effectively guide the learning of skeleton representation from the complete topology skeleton.
TABLE VIII: The influence of different skeleton encoders on NTU RGB+D 60 dataset with the joint stream.
Backbone Architecture NTU60
xsub xview
PCM3 GRU 74.1 (-10.3) 80.1 (-9.7)
UmURL Transformer 79.0 (-5.4) 84.7 (-5.1)
MS-G3D GCN 84.5 (+0.1) 89.9 (+0.1)
CTR-GCN GCN 84.5 (+0.1) 90.4 (+0.6)
ST-GCN (Our used) GCN 84.4 89.8
IV-E3 Influences of the Hyper-parameters
As shown in Fig. 6, we evaluate the influence of the hyper-parameters. For the hyper-parameter $\alpha$, we find that the teacher network should dominate the early learning period, and the progressive extent should be large for full exploration to align unpaired cross-modal data with high similarity. Meanwhile, the hyper-parameter $\beta$ exhibits a slight increase and decrease, showing that the hard alignment and the robustness of our proposed intra-modal similarity guidance should be required.
Figure 6: (a) The influence of progressive extent $\alpha$ at different ratios on the NTU RGB+D 60 dataset. (b) The influence of intra-modal similarity guidance extent $\beta$ at different ratios on the NTU RGB+D 60 dataset.
(a) Language (InfoNCE)
(b) Language (Ours)
(c) Vision (InfoNCE)
(d) Vision (Ours)
Figure 7: The t-SNE visualization of action representations obtained by original InfoNCE loss (7a, 7c) and our proposed loss(7b, 7d) with the different modality as the supervision on the NTU RGB+D 60 xsub dataset. 10 classes from the testing set are randomly selected for visualization, where points of the same color correspond to the same action category.
IV-F Qualitative Analysis
IV-F1 t-SNE Visualization of Representations
We employed t-SNE [63] on the NTU RGB+D 60 xsub benchmark to visualize the embedding distribution obtained by InfoNCE loss and our proposed loss with vision and language knowledge prompts as supervision, respectively. As shown in Fig. 7, we select 10 classes for visualization, where points of the same color correspond to the same action category. It reveals that language provides better guidance for representations of discriminative skeletons than vision knowledge. Moreover, the proposed soft alignment and progressive self-distillation strategies have the advantage of clustering similar representations, resulting in better inter-class compactness and intra-class separability.
IV-F2 Confusion Matrix
Here, we draw the confusion matrices of skeleton-based action recognition results with the different modalities as the supervision based on InfoNCE loss and our proposed strategies. As shown in Fig. 8, we can find that the proposed soft alignment strategies significantly improve results, especially with the vision knowledge prompts as the supervision. Compared to the InfoNCE loss, our proposed loss improves recognition accuracy by nearly 10% with the vision knowledge, demonstrating that aligning the high-similar unpaired data is necessary.
(a) Language (InfoNCE)
(b) Language (Ours)
(c) Vision (InfoNCE)
(d) Vision (Ours)
Figure 8: The confusion matrices of original InfoNCE loss (8a, 8c) and our proposed loss (8b, 8d) with the different modality as the supervision on the NTU RGB+D 60 xsub dataset under the linear probe evaluation.
V Conclusion
In this work, we propose a novel skeleton-based training paradigm with vision-language knowledge prompts as the supervision rather than extensive human-designed coarse-grained annotations. Leveraging large multimodal models, we generate the vision-language knowledge prompts to establish the vision-language action concept space, enriching the skeleton action space with fine-grained details. In addition, we design the intra-modal self-similarity and inter-modal cross-consistency softened targets to effectively control and guide the degree to which samples should be selectively pulled closer in the instance discrimination and knowledge distillation process. These noisy pairs come from the abovementioned action space consisting of vision-language prompts and their corresponding skeleton. During the inference phase, only the skeleton serves as the input for action recognition. Experimental results show that our method achieves superior performance and promises better skeleton action representations.

### v2

@@
 Abstract
-Supervised and self-supervised learning are two main training paradigms for skeleton-based human action recognition.
+Skeleton-based action representation learning aims to interpret and understand human behaviors by encoding the skeleton sequences, which can be categorized into two primary training paradigms: supervised learning and self-supervised learning.
 However, the former one-hot classification requires labor-intensive predefined action categories annotations, while the latter involves skeleton transformations (e.g., cropping) in the pretext tasks that may impair the skeleton structure.
@@
 Specifically, we establish the vision-language action concept space through vision-language knowledge prompts generated by pre-trained large multimodal models (LMMs), which enrich the fine-grained details that the skeleton action space lacks.
-Moreover, we propose the intra-modal self-similarity and inter-modal cross-consistency softened targets in the cross-modal contrastive process to progressively control and guide the degree of pulling vision-language knowledge prompts and corresponding skeletons closer.
+Moreover, we propose the intra-modal self-similarity and inter-modal cross-consistency softened targets in the cross-modal representation learning process to progressively control and guide the degree of pulling vision-language knowledge prompts and corresponding skeletons closer.
 These soft instance discrimination and self-knowledge distillation strategies contribute to the learning of better skeleton-based action representations from the noisy skeleton-vision-language pairs.
 During the inference phase, our method requires only the skeleton data as the input for action recognition and no longer for vision-language prompts.
-Extensive experiments show that our method achieves state-of-the-art results on NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD datasets.
-The code will be available in the future.
+Extensive experiments on NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD datasets demonstrate that our method outperforms the previous methods and achieves state-of-the-art results.
+Code is available at: https://github.com/cseeyangchen/C2VL.
 Index Terms:
-Action Recognition, Vision-Language, Cross-Modal, Self-Supervised, Contrastive Learning, Large Multimodal Models.
+Action Recognition, Vision-Language, Cross-Modal, Self-Supervised, Contrastive Learning, LMMs.
 I Introduction
 Action recognition has become a hot topic due to its wide range of application fields, such as human-computer interaction, sports analysis, and dangerous behavior warning [1], etc.
-In recent years, the continuous development of depth sensors like Kinect [2] has facilitated the easy acquisition of the human body’s skeleton (joints coordinate).
+In the past ten years, the continuous development of depth sensor cameras like Kinect [2] has facilitated the easy acquisition of the human body’s skeleton (3D joints coordinate).
 Compared with RGB-D sequences, the 3D skeleton is naturally robust to light, background noise, and viewpoint changes while also having the advantages of being lightweight and privacy-preserving.
@@
 On the other hand, self-supervised learning approaches generate positive-negative pairs by designing pretext tasks of data transformations (rotation, masking, and cropping, etc.) to explore task-agnostic action representation in a label-free manner based on contrastive learning [6, 7, 8].
-These methods leverage pretext tasks for unsupervised pre-training and then finetune for downstream tasks.
+These methods leverage pretext tasks for unsupervised pre-training and then fine-tuning them for downstream tasks.
 The results suggest that the self-supervised approaches have the potential to surpass supervised learning methods.
-However, while the abovementioned methods have proven effective in representing human actions, several challenges remain.
-(1) Obtaining numerous labeled data is extremely expensive and labor-intensive for supervised learning.
-Additionally, the reliance on one-hot vector classification restricts the generality of these methods since additional labeled data is needed to address any other tasks.
-(2) For self-supervised learning, skeleton data transformations (rotation, masking, cropping, etc.) of the pretext task may impair the structural information inherent in skeleton action, leading to information bias.
-Inspired by the recent success in multi-modal contrastive learning [9, 10, 11], we consider designing a novel skeleton training paradigm guided by vision-language knowledge to obviate numerous label annotations or compromise the integrity of the skeleton structures, as shown in Fig.
+However, while the aforementioned methods have proven effective in representing human actions, several challenges remain.
+(1) The one-hot vector classification in supervised learning requires time-consuming and labor-intensive annotating of numerous labels while restricting the generalization performance of these methods.
+(2) For self-supervised learning, data transformations (rotation, masking, cropping, etc.) of the pretext task may impair the structural information inherent in skeleton action, leading to information bias.
+For example, the motion of the ”hand rising” action will be impaired if the masking transformation is applied to the hand joints [8].
+Inspired by the recent success in multi-modal contrastive learning [9, 10, 11], we consider designing a novel skeleton training paradigm guided by cross-modal representation learning with auto-generated vision-language knowledge prompts, which can boost the task-agnostic generalization capacity while obviating numerous human-effort annotation and irreversible skeleton impairments, as shown in Fig.
 1.
 Specifically, our approach leverages sample-specific vision-language knowledge prompts as the supervision to guide the skeleton encoder in learning task-agnostic action representation, which is called C2VL.
-Firstly, we utilize LMMs (Grounding DINO [12] and LLaVA [13]) as the vision-language engines, employing text prompts and visual questions to generate one-to-one vision knowledge prompts related to human action images and one-to-one language knowledge prompts corresponding to action descriptions, respectively.
+Firstly, we utilize two LMMs (Grounding DINO [12] and LLaVA [13]) as the vision and language engines, respectively.
+The former is an open-set object detector, and the latter is a foundation language model.
+Then, we employ text prompts and visual questions for them to auto-generate one-to-one vision knowledge prompts related to human action images and one-to-one language knowledge prompts corresponding to action descriptions, respectively.
 Subsequently, the vision-language encoders extract features from these knowledge prompts to establish the vision-language action concept space, enriching object-related details and fine-grained descriptions absent in the skeleton action space.
@@
 In this section, we mainly introduce the related work of skeleton-based action recognition, multi-modal representation learning, and vision-language pretraining with noisy data.
-II-A Skeleton-based Action Recognition
+II-A Skeleton-based Action Recognition Methods
 Supervised skeleton-based action recognition has been the dominant paradigm.
@@
 They reconstruct skeleton structures by masked modeling [19, 20, 21, 22] or generate positive-negative pairs [8, 7, 6, 23, 24] through skeleton transformation (e.g., rotation, cropping) respectively to explore action representation.
-Unfortunately, these pretext tasks impair the topological structure of the skeleton, leading to information bias.
+Unfortunately, these pretext tasks impair the topological structure of the skeleton, leading to information bias [8].
 [Summary]: In this paper, we first explore the possibility of learning action representation for skeleton-based action recognition from other modalities (vision and language) without label annotation or skeleton damage.
-II-B Multimodal Representation Learning
-Numerous multimodal approaches have demonstrated the capacity to learn powerful representations from different modalities.
+II-B Multimodal Representation Learning Methods
+Numerous multimodal approaches have demonstrated the capacity to learn powerful representations from different modalities [25].
 Inspired by their success, foundational vision-language models such as CLIP [9] and ALIGN [26] have gained widespread adoption in the domain of RGB-based action recognition [27, 1, 28, 10, 29, 30, 31, 32, 11].
 Many approaches rely on human-designed coarse-grained category names for video action classification [10], pose estimation [29] and action generation [30], etc.
+PalmCohashNet utilizes multimodal co-learning hash technology for identity authentication [25].
 Much recent research gradually employs LLMs (GPT-3, etc.) to generate detailed action descriptions of category names for RGB-based action classification [31] and generation [32].
@@
 Second and more importantly, we utilize these one-to-one vision-language knowledge prompts as the supervisory signal to learn skeleton representation in a self-supervised learning paradigm.
-II-C Contrastive Learning with Noisy Pairs
+II-C Contrastive Learning with Noisy Pairs Methods
 Contrastive learning is the core of the recent multimodal self-supervised learning methods, which pulls the positive pairs closer and pushes the negative pairs away [33, 34].
@@
 III-A Vision-Language Action Concept Space
-For a given skeleton dataset $D\in\left\{S_{i}\right\}_{i=1}^{\left|\mathrm{D}\right|}$, we first generate the sample-level vision-language knowledge prompts $\left\{(V_{i},L_{i})\right\}_{i=1}^{\left|\mathrm{D}\right|}$ by offline LMMs (Grounding DINO [12] and LLaVA [13]).
+For a given skeleton dataset $D\in\left\{S_{i}\right\}_{i=1}^{N}$, we first generate the sample-level vision-language knowledge prompts $\left\{(V_{i},L_{i})\right\}_{i=1}^{N}$ by offline LMMs (Grounding DINO [12] and LLaVA [13]), where $N$ denotes the number of samples in dataset.
+$S_{i}$ is the $i$-th skeleton sample, $V_{i}$ is the $i$-th vision sample, and $L_{i}$ is the $i$-th language sample.
 These sample-level vision-language knowledge prompts have a one-to-one corresponding with skeleton sequence, which establishes the vision-language action concept space as shown in Fig.
@@
 These raw videos and skeleton sequences have one-to-one correspondence as they are captured simultaneously by Kinect devices.
-However, raw videos are challenged to capture fine-grained details and subtle movement changes because of interference from irrelevant background environments.
+However, raw videos are challenged to capture fine-grained details and subtle movements because of interference from irrelevant background environments.
 To mitigate this issue, we leverage a pre-trained open-set object detector called Grounding DINO [12], with text prompts “person” to crop the human action images as vision knowledge prompts precisely.
@@
 Answer the questions concisely” to comprehend and generate action descriptions to the vision action images autonomously.
-These language knowledge prompts provide sample-specific fine-grained detail descriptions in the language action concept space.
+These language knowledge prompts provide sample-specific fine-grained descriptions in the language action concept space.
 Figure 3: The establishment of vision-language action concept space.
@@
 7c.
+Figure 4: Histogram of similarity scores for positive pairs between skeleton and vision-language knowledge representations in cross-modal space after training with original InfoNCE loss.
+Figure 5: Histogram of similarity scores for negative pairs between skeleton and vision-language knowledge representations in cross-modal space after training with original InfoNCE loss.
+The red ellipse highlights the movement of arms.
 III-C Cross-modal Alignment with Soft Targets
@@
 where $\mathrm{\mathbf{J}}$ is an all-one matrix.
-The first term of Equation 7-10 aggregates the information from its cross opposite modality for dual-direction alignment, i.e., alignment from vision $\mathrm{\mathbf{v}}_{i}$ to skeleton $\mathrm{\mathbf{s}}_{j}$ is based on the probability that skeleton $\mathrm{\mathbf{s}}_{i}$ matched with vision $\mathrm{\mathbf{v}}_{j}$.
+The first term of Equation (7)-(10) aggregates the information from its cross opposite modality for dual-direction alignment, i.e., alignment from vision $\mathrm{\mathbf{v}}_{i}$ to skeleton $\mathrm{\mathbf{s}}_{j}$ is based on the probability that skeleton $\mathrm{\mathbf{s}}_{i}$ matched with vision $\mathrm{\mathbf{v}}_{j}$.
 Moreover, the second and third terms are designed to avoid positive pairs that are not sound correspondences for boosting negative pairs with high semantic similarity in cross-modalities.
-Figure 4: Histogram of similarity scores for positive pairs between skeleton and vision-language knowledge representations in cross-modal space after training with original infoNCE loss.
-Figure 5: Histogram of similarity scores for negative pairs between skeleton and vision-language knowledge representations in cross-modal space after training with original infoNCE loss.
-The red ellipse highlights the movement of arms.
 III-D Progressive Self-distillation Training Objectives
@@
 Specifically, we partition a batch of pairs into two subsets with the refreshed hyper-parameter $\alpha\in[0,1]$ at each epoch.
-The intra-modal self-similarity soft target is utilized to repair the first subset data $\mathrm{B_{intra}}=\left\lfloor\alpha\mathrm{B}\right\rfloor$ for estimation of the many-to-many relationships, while the inter-class modal cross-consistency is employed to guide the second subset data $\mathrm{B_{inter}}=\mathrm{B}-\left\lfloor\alpha\mathrm{B}\right\rfloor$ for boosting negative pairs with high semantic similarity.
+The intra-modal self-similarity soft target is utilized to repair the first subset data $\mathrm{B_{intra}}=\left\lfloor\alpha\mathrm{B}\right\rfloor$ for estimation of the many-to-many relationships, while the inter-class modal cross-consistency is employed to guide the second subset $\mathrm{B_{inter}}=\mathrm{B}-\left\lfloor\alpha\mathrm{B}\right\rfloor$ for boosting negative pairs with high semantic similarity.
 At the initial training phase, the value of $\alpha$ should be large for training stability.
@@
 Meanwhile, $\underline{S},\underline{V},\underline{L}\in\mathbb{R}^{\mathrm{B_{inter}}\times d}$ denotes the last $\mathrm{B_{inter}}$ rows of respective embeddings.
-In the inference phase, only the skeleton sequences and encoder $E_{s}$ are used for skeleton-based action recognition. The vision-language prompts and respective encoders will be dropped.
+In the inference phase, only the skeleton and encoder $E_{s}$ are used for skeleton-based action recognition. The vision-language prompts and respective encoders will be dropped.
 IV Experiments
@@
 In this experiment, we adopt the cross-subject benchmark following the previous studies.
-TABLE I: Comparison to the state-of-the-art methods for action recognition on the NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD II datasets under the linear evaluation protocol.
+TABLE I: Comparison to the state-of-the-art methods for action recognition accuracy on the NTU RGB+D 60, NTU RGB+D 120, and PKU-MMD II datasets under the linear evaluation protocol.
 The best and second-best results are highlighted in bold and underlined, respectively.
@@
 In the inference stage, we utilize multiple evaluation protocols, including the linear evaluation protocol, finetune protocol, semi-supervised evaluation protocol(1%, 5% or 10% labeled data), KNN evaluation protocol, and transfer learning evaluation protocol for action recognition.
+Meanwhile, following the previous studies [40, 47, 48, 49, 8, 50, 23, 51, 52, 24], several tasks (cross-view, cross-subject, and cross-setup) with different official training-testing split strategies are used to comprehensively validate the algorithm, preventing overfitting in the testing set.
 IV-C Comparison to the State-of-the-art
@@
 Xsub Xview
-PCM3 GRU 74.1 (-10.3) 80.1 (-9.7)
+ST-GCN (Our used) GCN 84.4 89.8
+[…]
 UmURL Transformer 79.0 (-5.4) 84.7 (-5.1)
@@
 CTR-GCN GCN 84.5 (+0.1) 90.4 (+0.6)
-ST-GCN (Our used) GCN 84.4 89.8
 IV-E3 Influences of the Hyper-parameters